The govemment starts a new passenger train service linking two large cities. After two years operating costs have increased. To compensate for this the government introduces a ticket levy of S50 per passenger. Your task is to determine the consequences of this levy. The supply and demand equations are given by Demand Equation p 1800 9aq Supply Equation: p-0.0000592 3ba + 150 where q is the mumber of passengers (in thousands per year) ; p is the train ticket price in dollars. a and b are coefficients to be chosen below. Steps to complete to undertake the investigation 1. Obtain two random numbers for a and b between 0.1 and 0.9 inclusive. Use the Excel random function-Rand0 or the RANDOM key on your calculator 2. Find the equilibrium price and quantity algebraically, without the proposed ticket levy 3. Update the demand equation to include the S50 surcharge. 4. Use algebra to find the new equilibrium price and quantity with the levy 5. Use Excel to graph on the same set of axes, the original supply and demand equations (from part 1) and the updated demand equation with the proposed levy (part 3). 6. Use your results to complete the following summary table. Round answers to two decimal Before l After l Difference Equilibrium quantity Equilibrium price: consumer pays Equilibrium price before l Revenue government receives (without l Revenue government receives (with levy) LEAVE THIS CELL BLANK
Answer #1

Let a=b=1/3

so,

WITHOUT LEVY

demand: P = 1800 - 3Q

Supply : P = 0.00005 Q2 + Q + 150

WITH LEVY

demand: P + 50 = 1800 - 3Q

Supply : P = 0.00005 Q2 + Q + 150

Before Levy After levy Difference
Eq. quantity 410 398 12
Eq. price: consumer pay 570 556 4
Eq. price: before levy 568.4 555.9 12.5
govt Revenue(without levy) 1.6 0.1 1.5
govt Revenue(with levy) -